OpenStudy (anonymous):

Someone please help me! Fan and medal (:

OpenStudy (anonymous):

Here is the question.^^^^

OpenStudy (anonymous):

Find the value of the variables.

OpenStudy (anonymous):

Cos1= .54 cos45= .52 7/14 = .5 = 1/2

OpenStudy (anonymous):

$$ cos(x) = \frac{7}{14} = \frac{\cancel 7}{\cancel 7 \cdot 2} = \frac{1}{2} \\ x = \pm60^\circ + 360^\circ K = \pm \frac{\pi}{3} + 2\pi K $$